Afternoon tasks: ---------------- DA and instrument checkout completed with no problems. Jon B. & Wendell worked some more this afternoon on the secondary galil motor connections. They replaced pins and replaced spliced wires. Steph ran the secondary back & forth several times, but could not reproduce the "not stopping on a full step error". However, during opening, we again saw the secondary E Galil failure to find a full step. We homed the actuator, noting an error of 148 microsteps. When we re-enabled the collimation process, we saw the error again almost immediately. We homed again, this time seeing an error of 1 microstep. We did not see the error again. Robert & Steph verified that the same fiducial tables were used for both pointing models taken this dark run. The difference between the two models that was noted last night is not due to the fiducial tables. Pete & Craig called Russell to see if there were any TCC related problems that could have caused the difference. We eliminated axisDtime errors and problems with the earth orientation data block as possible causes. The mystery continues. We opened at 20:30 under partly cloudy skies. Cart 1, no star with either model. Cart 3. field at ra=313, dec=-6. fk5infiber went to 260 +47? second fk5infiber went to nearby star. Offset ra+12, dec+2 to center star, then we got guide stars. Spectro 1 fill faults at 22:19 did not correspond to actual warming. Closed by high humdity at 00:05 MDT. i+trim2 reported -6.68V for CCD61. Watched seemed to report this about midnight (while we were closing), but a plot showed this had been the value since 10am this morning.